OpenStudy (anonymous):

A negatively charged particle with a mass of 6.00x10^-8kg is travelling due east at 75.0m/s; it then enters a 0.550T uniform magnetic field. The particle moves through three-quarters of a circle in 4.50x10-3s, whereupon it leaves the field heading due south. All during the motion the particle moves perpendicular to the magnetic field. a. what is the magnitude of the magnetic force on the particle? b. Which direction does the magnetic field point? c. What is the particle's charge?

OpenStudy (anonymous):

\[F=q v \times B\] also from given mass move in anti clock direction direction of magnetic field is down-word c. \[m v^2 /r =qv \times B\] you can calculate r by \[\frac{ 3 }{ 4 } 2 \pi r= vt\]

OpenStudy (anonymous):

Alright thanks :D I hope you're right!

OpenStudy (anonymous):

oh wait charge particle is negative so direction of magnetic force would be upward not downward.